when the smaller of two consecutive intergers is added to four times the larger the result is 49
Answer by algebrahouse.com(1659)

x = first consecutive integer
x + 1 = second consecutive integer {consecutive integers increase by 1}

x + 4(x + 1) = 49 {the smaller is added to 4 times the larger to equal 49}
x + 4x + 4 = 49 {used distributive property}
5x + 4 = 49 {combined like terms}
5x = 45 {subtracted 4 form each side}
x = 9 {divided each side by 5}
x + 1 = 10 {substituted 9, in for x, into x + 1}

9 and 10 are the two consecutive integers

For more help from me, visit: www.algebrahouse.com